The displacement of an object executing simple harmonic oscillations is given by:


x = 0.02 sin 2π ( t + 0.01 ) m


Determine (a) amplitude of the oscillatory motion, (b) time period of oscillation , (c) maximum velocity, (d) maximum acceleration, and (e) initial displacement of the object.

1
Expert's answer
2022-03-13T18:49:30-0400

Solution.

"x=0.02sin2\pi(t+0.01)=0.02sin(2\pi t+0.02\pi)m;"

"x=Asin(\omega t+\phi_0);"

"a)A=0.02m;"

"b)\omega=\dfrac{2\pi}{T}\implies T=\dfrac{2\pi}{\omega};" "\omega=2\pi s;"

"T=\dfrac{2\pi}{2\pi}=1s;"

"c)v_{max}=A\omega;"

"v_{max}=0.02\sdot2\sdot3.14=0.1256m/s;"

"d) a_{max}=A\omega^2;"

"a_{max}=0.02\sdot(2\sdot 3.14)^2=0.789 m/s^2;"

"e)x_0=0.02sin0.02\pi=0.0011m;"

Answer: "a)A=0.02m;"

"b)T=1s;"

"c)v_{max}=0.1256m\/s;"

"d)a_{max}=0.789 m\/s^2;"

"e)x_0=0.0011m."